OpenStudy (anonymous):

Find an equation in standard form for the hyperbola with vertices at (0, ±3) and foci at (0, ±7)

OpenStudy (jdoe0001):

so, I gather you know the equation of a hyperbola?

OpenStudy (jdoe0001):

well @n14r96 the vertices are at (0, +3) and (0, -3) the "x" doesn't change the "y" does meaning is going upwards, thus is a vertical traversal axis

OpenStudy (jdoe0001):

the distance from one vertex to the other is 3+3, so 6 half-way through is the (h,k) center the distance for the "a" element in the formula is the distance from the center to the vertex so the vertices are at (0, +3) and (0, -3), so the "a" distance is 3 so a = 3 the foci are at (0, ±7) the distance from the center to a focus is "c" you center is at (0,0) your foci at (0, ±7) so the "c" distance is 7 c = 7 so, what about "b"? \(c^2=a^2+b^2 \implies 7^2 = 3^2+b^2\) solve for \(b^2\) the equation uses \(a^2 \ and \ b^2\) so since a = 3, \(a^2 = 9\)

OpenStudy (jdoe0001):

and of course, your center is at (0,0)

OpenStudy (anonymous):

so y^2/49-x^2/9=1 ???